# post-last exit time process and its application to Loss-Given-Default Distribution

### v3

@@ 2.1. Elements of Linear Diffusion
with the Wronskian $w_{q}:=\psi_{q}^{+}(x)\phi_{q}(x)-\psi_{q}(x)\phi_{q}^{+}(x)=\psi_{q}^{-}(x)\phi_{q}(x)-\psi_{q}(x)\phi_{q}^{-}(x)$. We have $G_{q}(x,y)=\int_{0}^{\infty}e^{-qt}p(t;x,y){\rm d}t$ for $x,y\in\mathcal{I}$.
@@
2.2. Post-last exit time process
@@
We are interested in the distribution of a transient diffusion $X$ after the last exit time from $\alpha$; i.e., $\mathbb{P}_{x}(X_{t}\in{\rm d}y,L_{\alpha}<t)$. Recall that for transient diffusion $L_{\alpha}<\infty$ a.s. Let us define
$h_{\alpha}(x):=\mathbb{P}_{x}(L_{\alpha}=0)$
@@
Under Assumption 1, the transition density (with respect to the speed measure $m({\rm d}y)$) of the post-last exit time process $(X_{t})_{t>L_{\alpha}}$ is given by (2.5). More precisely,
$p^{h_{\alpha}}(t-u;z,y)=\frac{h_{\alpha}(y)}{h_{\alpha}(z)}p(t-u;z,y),\qquad y<z<\alpha,\quad u\in(L_{\alpha},t)$ (2.6)
for $u$ arbitrarily close to $L_{\alpha}$: $z$ is arbitrarily close to $\alpha$.
@@
Let us fix any two points $\𝛼x,y\in\mathcal{I}\backslash\{\alpha\}$ such that $y<\alpha$ and consider $\mathbb{P}_{x}(X_{t}\in{\rm d}y,L_{\alpha}<t)$ for an arbitrary $t>0$. By Bayes’ rule, we have
$\mathbb{P}_{x}(X_{t}\in{\rm d}y,L_{\alpha}<t)$ $=\mathbb{P}_{x}\left(L_{\alpha}<t\middle|X_{t}\in{\rm d}y\right)\mathbb{P}_{x}(X_{t}\in{\rm d}y)$
$=\mathbb{P}_{x}(L_{\alpha}\circ\theta_{t}=0|X_{t}\in{\rm d}y)\mathbb{P}_{x}(X_{t}\in{\rm d}y)$
$=\mathbb{P}_{y}(L_{\alpha}=0)p(t;x,y)m({\rm d}y)$ (2.7)
where the second line is due to (2.1) and the third line is due to the Markov property at time $t$. Our objective here is to identify the transition semigroup of the post-last exit time process and for this purpose, let us define
$T_{z}:=\inf\{s>L_{\alpha}:X_{s}=z\}.$ (2.8)
Then, we have
$\mathbb{P}_{x}(X_{t}\in{\rm d}y,L_{\alpha}<t)$ $=\mathbb{P}_{y}(L_{\alpha}=0)\mathbb{E}_{x}[p(t;x,y)\mbox{1}\mbox{l}_{\{T_{z}<t\}}+p(t;x,y)\mbox{1}\mbox{l}_{\{T_{z}\geq t\}}]m({\rm d}y)$
$=\mathbb{P}_{y}(L_{\alpha}=0)\mathbb{E}_{x}[p(t;x,y)\mbox{1}\mbox{l}_{\{T_{z}<t\}}]m({\rm d}y)+\mathbb{P}_{y}(L_{\alpha}=0)p(t;x,y)\mathbb{P}_{x}(T_{z}\geq t)m({\rm d}y).$
This result with (2.2) yields
$\mathbb{P}_{x}(X_{t}\in{\rm d}y,L_{\alpha}<t)\mathbb{P}_{x}(T_{z}<t)=\mathbb{P}_{y}(L_{\alpha}=0)\mathbb{E}_{x}[p(t;x,y)\mbox{1}\mbox{l}_{\{T_{z}<t\}}]m({\rm d}y).$ (2.9)
Now we rewrite (2.9) with a view from time $T_{z}$ to $t$ instead of that from time zero to $t$. Let us first undertake the second term of the right-hand side by observing that $p(t;x,y)=\int_{\mathcal{I}}p(s;x,u)p(t-s;u,y)m({\rm d}u)$ for $0<s<t$.
For an arbitrary random time $T<\infty$ a.s., define a random measure $\tilde{p}(T;u,v)$ such that $\mathbb{E}_{x}[\tilde{p}(T;u,v)m({\rm d}v)]=\mathbb{P}_{u}(X_{T}\in{\rm d}v)$ for $u,v\in\mathcal{I}$. Note that $\tilde{p}(T_{z};x,v)m({\rm d}v)=0$ a.s. when $v\neq z$. This is because by Fubini’s theorem
$E_{x}\left[\int_{\mathcal{I}}\tilde{p}(T_{z};x,v)m({\rm d}v)\right]$ $=\int_{\mathcal{I}}\mathbb{P}_{x}(X_{T_{z}}\in{\rm d}v)=\int_{\mathcal{I}}\int_{0}^{\infty}\mathbb{P}_{x}(X_{s}\in{\rm d}v,T_{z}\in{\rm d}s)$
$=\int_{0}^{\infty}\mathbb{P}_{x}(X_{s}\in{\rm d}z,T_{z}\in{\rm d}s)=\mathbb{P}_{x}(X_{T_{z}}\in{\rm d}z)=\mathbb{E}_{x}[\tilde{p}(T_{z};x,v)m({\rm d}v)\delta_{z}({\rm d}v)]$
and $\mathbb{P}_{x}(X_{T_{z}}\in{\rm d}z)=1$ by the definition of $T_{z}$. Hence the random measure $\tilde{p}(T_{z};x,v)m({\rm d}v)$ has a point mass at $z$ almost surely. Then, together with (2.5), the second term on the right-hand side of (2.9) becomes
@@
$\mathbb{E}_{x}[p(t;x,y)\mbox{1}\mbox{l}_{\{T_{z}<t\}}]m({\rm d}y)$ $=\mathbb{E}_{x}\left[\int_{0}^{t}\int_{\mathcal{I}}p(s;x,u)p(t-s;u,y)m({\rm d}u)\mbox{1}\mbox{l}_{\{T_{z}\in{\rm d}s\}}\right]m({\rm d}y)$
$=\mathbb{E}_{x}\left[\int_{\mathcal{I}}\tilde{p}(T_{z};x,u)\tilde{p}(t-T_{z};u,y)m({\rm d}u)\mbox{1}\mbox{l}_{\{T_{z}<t\}}\right]m({\rm d}y)$
$=\mathbb{E}_{x}\left[\tilde{p}(t-T_{z};X_{T_{z}},y)\mbox{1}\mbox{l}_{\{T_{z}<t\}}\right]m({\rm d}y)$
@@
$=\mathbb{E}_{x}\left[\frac{h_{\alpha}(X_{T_{z}})}{h_{\alpha}(y)}\tilde{p}^{h_{\alpha}}(t-T_{z};X_{T_{z}},y)\mbox{1}\mbox{l}_{\{T_{z}<t\}}\right]m({\rm d}y)$
@@
$=\mathbb{E}_{x}\left[\frac{1}{h_{\alpha}(y)}\tilde{p}^{h_{\alpha}}(t-T_{z};X_{T_{z}},y)\mbox{1}\mbox{l}_{\{T_{z}<t\}}\right]m({\rm d}y)$
where $\tilde{p}^{h_{\alpha}}(T;u,v)m({\rm d}v):=\frac{h_{\alpha}(v)}{h_{\alpha}(u)}\tilde{p}(T;u,v)m({\rm d}v)$ for an arbitrary random time $T<\infty$ a.s. and $u<\alpha$ similarly to (2.5). Note that in the last equality we used $h_{\alpha}(X_{T_{z}})=\mathbb{P}_{X_{T_{z}}}(L_{\alpha}=0)=1$ by the definition of $T_{z}$. This result simplifies (2.9) to
$\mathbb{P}_{x}(X_{t}\in{\rm d}y,L_{\alpha}<t)\mathbb{P}_{x}(T_{z}<t)$ $=\int_{0}^{t}p^{h_{\alpha}}(t-s;z^{*},y)\mathbb{P}_{x}(T_{z}\in{\rm d}s)m({\rm d}y)$ (2.10)
where we use $z^{*}$ in $p^{h_{\alpha}}(\cdot;\cdot,\cdot)$ rather than a mere $z$ to emphasize and recall that this is the point attained by $X$ at the time $T_{z}$ defined in (2.8): $z^{*}:=X_{T_{z}}=z$.
We have so far shown the following: in our process of rewriting the second term on the right-hand side of (2.9), we are able to use (2.5) to obtain (2.10). It follows that the remaining task is to rewrite the left-hand side of (2.10) with a view from time $T_{z}$. The corresponding expression is obtained by looking at the shifted path $(X_{u}\circ\theta_{T_{z}})(\omega)=X_{u}(\theta_{T_{z}}\omega)=X_{T_{z}+u}(\omega)$. Recall the shift operator in (2.2). We have
@@
$\mathbb{P}_{x}(X_{t}\in{\rm d}y,L_{\alpha}<t)$ $=\mathbb{P}_{x}(X_{t}\in{\rm d}y,T_{z}<t)=\mathbb{E}_{x}[\mathbb{E}_{x}(\mbox{1}\mbox{l}_{\{X_{t}\in{\rm d}y,T_{z}<t\}}|\mathcal{F}_{T_{z}})]$
$=\mathbb{E}_{x}[\mbox{1}\mbox{l}_{\{T_{z}<t\}}\mathbb{P}_{x}(X_{t-T_{z}}\circ\theta_{T_{z}}\in{\rm d}y|\mathcal{F}_{T_{z}})]=\mathbb{P}_{x}(X_{t-T_{z}}\circ\theta_{T_{z}}\in{\rm d}y,T_{z}<t).$ (2.11)
Finally, division by $\mathbb{P}_{x}(T_{z}<t)$ together with (2.2) transforms (2.10) into
@@
$\mathbb{P}_{x}(X_{t-T_{z}}\circ\theta_{T_{z}}\in{\rm d}y,T_{z}<t)$ $=\left(\int_{0}^{t}p^{h_{\alpha}}(t-s;z^{*},y)\frac{\mathbb{P}_{x}(T_{z}\in{\rm d}s,T_{z}<t)}{\mathbb{P}_{x}(T_{z}<t)}\right)m({\rm d}y)$
@@
$=\left(\int_{0}^{t}p^{h_{\alpha}}(t-s;z^{*},y)\mathbb{P}_{x}(T_{z}\in{\rm d}s|T_{z}<t)\right)m({\rm d}y)$ (2.12)
@@
where the randomness comes only from $T_{z}$. But again, by the path continuity, we can make $z$ as close to $\alpha$ as we please. Therefore, for any time $L_{\alpha}+s,s>0$, the post- last exit time process is governed by (2.6) for $s$ arbitrarily close to zero. ∎
@@
Although $h_{\alpha}(x)$ is not an excessive function, the transform (2.6) has the same form as if it were an excessive function. Therefore, we can compute the infinitesimal drift and diffusion parameters of the transform (2.5) in the same way as if it is an $h$-transform. But there is a caveat that $x\neq\alpha$.
@@
Suppose that $Y$ is a diffusion (does not have to be transient) and let $h$ be an excessive function. Let us denote the infinitesimal drift and diffusion parameters of the $h$-transform by $\mu^{h}(\cdot)$ and $\sigma^{h}(\cdot)$, respectively. The following result slightly generalizes the argument in Section 15.9 of Karlin and Taylor (1981).
@@
Assume $h$ is an excessive function and its derivative $h^{\prime}(y)$ exists for $y\in\mathcal{I}$. The infinitesimal drift $\mu^{h}(y)$ and $\sigma^{h}(y)$ of the $h$-transform of $Y$ are
[…]
@@
Turning to our case,
$h_{\alpha}(x)=\mathbb{P}_{x}(L_{\alpha}=0)=\frac{s(\alpha)-s(x)}{s(\alpha)-s(\ell)},\quad x\leq\alpha.$ (2.14)
@@
The point $\alpha$ is an entrance boundary for the post-exit time process $X^{h_{\alpha}}_{t}(\omega):=\{\omega(t):L_{\alpha}<t<H_{\ell},\mathbb{P}^{h_{\alpha}}_{\alpha}\}$.

### v4

@@ 2.3. Asset-to-debt ratio process
We have the normalized process, with state space $\mathcal{I}=(\ell,r)=(-\infty,+\infty)$,
@@
by setting $M:=\frac{\mu-\frac{1}{2}\sigma^{2}-r}{\sigma}$. Now let us introduce the post-last passage time process
$Z_{t}:=\left(\frac{1}{\sigma}\ln(Y_{t})\right)\circ\theta(\lambda_{\alpha}),\quad t>0.$
We need first to identify the dynamics of the process $Z$ to estimate its value at the default time $\xi$, which we shall undertake in Section 3.2 after first providing new mathematical results in the next section.
3. Two mathematical results concerning last passage time
In this section, let us insert two mathematical results that apply to general transient linear diffusions. Both are, we believe, interesting in their own rights (since the results are general) and useful for the purpose of credit risk modeling. Due to the fact that our proposed modeling handles a whole lifetime of diffusion (the asset-to-debt ratio process), pre- and post- last passage time state can be consistently analyzed with these mathematical results.
3.1. Occupancy time above $\alpha$ before the last passage time
The first result is concerned with the occupation time above a certain level, say $\alpha\in\mathcal{I}=(\ell,r)$, before transient diffusion is killed. There is a well-known example in the classic textbook of Karlin and Taylor (1981) (see its Section 8-B, Chapter 15) where a diffusion is killed at its first passage to the left boundary and the time change by the occupation time of the positive axis is considered. The last passage time $\lambda_{\alpha}$ represents the final instance when the process is in the area above level $\alpha$. We provide the distribution, in terms of its Laplace transform, of the occupancy time above $\alpha$ for a general transient diffusion (A recurrent diffusion can be made transient if we set some finite state as a killing boundary). To our knowledge, this distribution is not available in the literature.
In the context of credit risk, this translates into how long the company’s asset-to-debt ratio stays above level $\alpha$. We can compute the distribution just by using estimated parameters in (2.5). This information should be of value to the management, banks, and investors. A ready-to-go practice is to set the killing boundary $\ell=0.4$ for the process $\frac{1}{\sigma}{\rm d}\ln(Y_{t})$, assuming that its LGD is around the market convention of $60\%$.
Fix some $\alpha\in\mathcal{I}$ as usual. Consider an occupation time of the region above $\alpha$
$\Gamma_{+}(t):=\int_{0}^{t}\mathbf{1}_{\{X_{s}\geq\alpha\}}{\rm d}s,\quad t\geq 0,$
together with its right inverse:
$\Gamma_{+}^{-1}(t):=\inf\{s:\Gamma_{+}(s)>t\}.$
Now let us define $\lambda^{A}_{\alpha}:=\Gamma_{+}(\lambda_{\alpha})$, which is the occupation time of $X$ above level $\alpha$ up to the last passage time $\lambda_{\alpha}$. Under Assumption 1, it is almost sure finite. We will use the following time-changed process:
$\hat{X}^{A}(t):=X(\Gamma_{+}^{-1}(t)).$
This process $\hat{X}^{A}$ is illustrated as FIG. 4 in the aforementioned example of Karlin and Taylor (1981) and the letter $\zeta$ there should read $\lambda^{A}_{\alpha}$ in the current notation.
Let us introduce another diffusion $X^{A}$ on $\alpha,r)$ for which $\alpha$ is a reflecting boundary. The process $X^{A}$ has the same speed measure and scale function as $\hat{X}^{A}$ (hence the same as $X$ as shown in Dynkin (1965, Theorem 10.12)) but its killing measure is zero. To be precise, we denote
$\hat{X}^{A}(t)=\begin{cases}X^{A}(t),\quad 0\leq t<\lambda_{\alpha}^{A}\\ \Delta,\quad t\geq\lambda_{\alpha}^{A}\end{cases}$
to distinguish the killed process $\hat{X}^{A}$ from $X^{A}$.
Before we proceed further, we need to introduce the local time at $\alpha$ for $\hat{X}^{A}$ by denoting
$\hat{L}^{A}(t):=\hat{L}^{A}(t,\alpha)=L(\Gamma^{-1}_{+}(t),\alpha)$
where $L(\cdot,\alpha)$ is the local time of $X$ at $\alpha$ which we denote by $L(\cdot)$ for brevity. We note that $X^{A}(t)=\hat{X}^{A}(t)$ and $L^{A}(t)=\hat{L}^{A}(t)$ on $0,\lambda_{\alpha}^{A}]$ where $L^{A}(t)$ is the local time at $\alpha$ for $X^{A}$. Due to this fact, we can deal with $L^{A}(\cdot)$ rather than $\hat{L}^{A}(\cdot)$ in the proof of the next proposition.
Let us also define the inverse local time process $\rho^{A}(s):=\inf\{t:L^{A}(t)>s\}$ for $X^{A}$ (see (Getoor, 1979, Section 7)) and also the q-Green function of $X^{A}$, denoted by $G_{q}^{A}$. The latter is, after some computation, $G_{q}^{A}(\alpha,\alpha)=\frac{\phi_{q}(\alpha)}{-\phi_{q}^{+}(\alpha)}$ where $\phi_{q}$ is in (2.2) and $\phi_{q}^{+}$ is the right derivative with respect to the scale function $s$.
Proposition 3.1.
Suppose that the state space of a general transient linear diffusion $X$ is $\mathcal{I}=(\ell,r)$. Under Assumption 1, the Laplace transform of $\lambda^{A}_{\alpha}$, the occupation time above level $\alpha$ before the last passage time $\lambda_{\alpha}$, is
$\mathbb{E}_{x}\left[e^{-q\lambda^{A}_{\alpha}}\right]=\frac{G^{A}_{q}(x,\alpha)}{G^{A}_{q}(\alpha,\alpha)+G(\alpha,\alpha)},\quad q>0,\quad x\geq\alpha,$ (3.1)
where $G(\alpha,\alpha)=s(\alpha)-s(\ell)$ (see (2.1)). Note that $\ell$ can be a finite ($\alpha>\ell$) number or negative infinity.
Proof.
Let $\mathbf{e_{q}}$ be an exponential random variable, independent of $X$, with rate $q>0$. By the independence of $X$ and $\mathbf{e_{q}}$, we have
$\mathbb{E}_{\alpha}\left[e^{-q\lambda^{A}_{\alpha}}\right]=\mathbb{P}_{\alpha}(\lambda^{A}_{\alpha}<\mathbf{e_{q}})=\mathbb{P}_{\alpha}(L^{A}(\lambda^{A}_{\alpha})<L^{A}(\mathbf{e_{q}})).$
Since $\lambda^{A}_{\alpha}:=\Gamma_{+}(\lambda_{\alpha})$ and $L^{A}(\lambda^{A}_{\alpha})=L(\Gamma^{-1}_{+}(\lambda^{A}_{\alpha}))$, we have
$L^{A}(\lambda^{A}_{\alpha})=L(\Gamma^{-1}_{+}(\Gamma_{+}(\lambda_{\alpha}))=\lim_{t\rightarrow\lambda_{\alpha}-}L(t)=L(\lambda_{\alpha}).$
By the argument in Section A.1, $L(\lambda_{\alpha})<\infty$ is the local time when the first excursion (from $\alpha$) of infinite length occurs and is exponentially distributed, as a Poisson arrival time, with rate $\frac{1}{G(\alpha,\alpha)}$. In other words, corresponding to local time $L(\lambda_{\alpha})$, we have a point (of the Poisson point process) representing that length of excursion.
On the other hand, it is known that the random variable $L^{A}(\mathbf{e_{q}})$ is exponentially distributed and
$\mathbb{P}_{\alpha}(L^{A}(\mathbf{e_{q}})>s)=\mathbb{P}_{\alpha}(\rho^{A}(s)<\mathbf{e_{q}})=\mathbb{E}_{\alpha}\left[e^{-q\rho^{A}(s)}\right]=\exp\left(-\frac{s}{G^{A}_{q}(\alpha,\alpha)}\right).$
The random variable $L(\lambda_{\alpha})$ is independent of any occasion that corresponds to a certain length of excursion because of the Markov property employed at each time an excursion begins. In particular, it is independent of $L^{A}(\mathbf{e_{q}})$. Therefore, we can compute
$\mathbb{P}_{\alpha}(L^{A}(\lambda^{A}_{\alpha})<L^{A}(\mathbf{e_{q}}))=\frac{\frac{1}{G(\alpha,\alpha)}}{\frac{1}{G(\alpha,\alpha)}+\frac{1}{G^{A}_{q}(\alpha,\alpha)}},$
which provides (3.1) with $x=\alpha$. For a general $x>\alpha$, by the strong Markov property at $H_{\alpha}$, we compute
$\mathbb{E}_{x}[e^{-q\lambda_{\alpha}}]=\mathbb{E}_{x}[e^{-q(H_{\alpha}+\lambda_{\alpha}\circ\theta_{H_{\alpha}})}]=\mathbb{E}_{x}[e^{-qH_{\alpha}}]\cdot\mathbb{E}_{\alpha}[e^{-q\lambda_{\alpha}}]$
and use $\mathbb{E}_{x}[e^{-qH_{\alpha}}]=\frac{\phi_{q}(x)}{\phi_{q}(\alpha)}$ (see (2.2)) and (2.3) to obtain (3.1). Note that $\phi_{q}^{A}(x)=\phi_{q}(x)$ on $\alpha,r)$ where $\phi_{q}^{A}$ denotes the decreasing solution of $\mathcal{G}^{A}f=qf$ with the generator $\mathcal{G}^{A}$ of $X^{A}$. ∎
For a comment on computations of $G_{q}^{A}(\cdot,\cdot)$ in various cases, see Appendix A.2.
3.2. Post-last passage time process
In this subsection, we shall identify a general form of the transition density of post-last passage time process in the class of general transient linear diffusions.
@@ 3.2. Post-last passage time process
We are interested in the distribution of a general transient diffusion $X$ after the last passage time from $\alpha$; i.e., $\mathbb{P}_{x}(X_{t}\in{\rm d}y,\lambda_{\alpha}<t)$. Recall that for transient diffusion $\lambda_{\alpha}<\infty$ a.s. Let us define
$h_{\alpha}(x):=\mathbb{P}_{x}(\lambda_{\alpha}=0)$ (3.2)
@@
Suppose that the state space of a general transient linear diffusion $X$ is $\mathcal{I}=(\ell,r)$. Under Assumption 1, the transition density (with respect to the speed measure $m({\rm d}y)$) of the post-last passage time process $(X_{t})_{t>\lambda_{\alpha}}$ is given by (3.3). More precisely,
$p^{h_{\alpha}}(t-u;z,y)=\frac{h_{\alpha}(y)}{h_{\alpha}(z)}p(t-u;z,y),\qquad y<z<\alpha,\quad u\in(\lambda_{\alpha},t)$ (3.4)
for $u$ arbitrarily close to $\lambda_{\alpha}$: $z$ is arbitrarily close to $\alpha$.
@@
[…]
$\mathbb{P}_{x}(X_{t}\in{\rm d}y,\lambda_{\alpha}<t)$ $=\mathbb{P}_{x}(\lambda_{\alpha}<t\mid X_{t}\in{\rm d}y)\cdot\mathbb{P}_{x}(X_{t}\in{\rm d}y)=\mathbb{P}_{x}(\lambda_{\alpha}\circ\theta_{t}=0|X_{t}\in{\rm d}y)\cdot\mathbb{P}_{x}(X_{t}\in{\rm d}y,H_{z}<t)$
@@
$=\mathbb{P}_{y}(\lambda_{\alpha}=0)\cdot\mathbb{E}_{x}\left[\mathbb{E}[\mbox{1}\hskip-2.5pt\mbox{l}_{\{X_{t}\in{\rm d}y\}}\cdot\mbox{1}\hskip-2.5pt\mbox{l}_{\{H_{z}<t\}}\mid\mathcal{F}_{H_{z}}]\right]$
$=h_{\alpha}(y)\cdot\mathbb{E}_{x}\left[p(t-H_{z};X_{H_{z}},y)m({\rm d}y)\cdot\mbox{1}\hskip-2.5pt\mbox{l}_{\{H_{z}<t\}}\right]$
@@
$=h_{\alpha}(y)\cdot\mathbb{E}_{x}\left[\left(p(t-H_{z};X_{H_{z}},y)\mbox{1}\hskip-2.5pt\mbox{l}_{\{\lambda_{\alpha}<H_{z}<t\}}+p(t-H_{z};X_{H_{z}},y)\mbox{1}\hskip-2.5pt\mbox{l}_{\{H_{z}<\lambda_{\alpha}<t\}}\right)m({\rm d}y)\right]$
where we used Markov property in the second and strong Markov property together with (3.2) in the third line. Note that $\mathbb{E}_{x}\left[h_{\alpha}(y)\cdot p(t-H_{z};X_{H_{z}},y)\cdot\mbox{1}\hskip-2.5pt\mbox{l}_{\{\lambda_{\alpha}>t>H_{z}\}}\right]=\mathbb{E}_{x}\left[\mathbb{E}_{x}\left[\mathbb{P}_{y}(\lambda_{\alpha}=0)\cdot\mbox{1}\hskip-2.5pt\mbox{l}_{\{X_{t}\in{\rm d}y\}}\cdot\mbox{1}\hskip-2.5pt\mbox{l}_{\{H_{z}<t\}}\mid\mathcal{F}_{H_{z}}\right]\cdot\mbox{1}\hskip-2.5pt\mbox{l}_{\{\lambda_{\alpha}>t\}}\right]=0.$ Let us examine separately the two terms on the right-hand side. The first term implies that $H_{z}$ occurs after $\lambda_{\alpha}$. Hence we use the transform (3.3) to obtain
@@
$\mathbb{E}_{x}\left[p(t-H_{z};X_{H_{z}},y)\mbox{1}\hskip-2.5pt\mbox{l}_{\{\lambda_{\alpha}<H_{z}<t\}}\right]$ $=\mathbb{E}_{x}\left[\frac{h_{\alpha}(X_{H_{z}})}{h_{\alpha}(y)}p^{h_{\alpha}}(t-H_{z};X_{H_{z}},y)\mbox{1}\hskip-2.5pt\mbox{l}_{\{\lambda_{\alpha}<H_{z}<t\}}\right]=\mathbb{E}_{x}\left[\frac{1}{h_{\alpha}(y)}p^{h_{\alpha}}(t-T_{z};X_{T_{z}},y)\mbox{1}\hskip-2.5pt\mbox{l}_{\{\lambda_{\alpha}<H_{z}<t\}}\right]$
where the last equality is due to $h_{\alpha}(X_{H_{z}})=\mathbb{P}_{X_{H_{z}}}(\lambda_{\alpha}=0)=1$ and $H_{z}=T_{z}$ on the set $\{\lambda_{\alpha}<H_{z}<t\}$.
To deal with the second term, we define, for an arbitrary random time $T<\infty$ a.s., a random measure $\tilde{p}(T;u,v)$ such that $\mathbb{E}_{x}[\tilde{p}(T;u,v)m({\rm d}v)]=\mathbb{P}_{u}(X_{T}\in{\rm d}v)$ for $u,v\in\mathcal{I}$. Note that $\tilde{p}(T_{z};x,v)m({\rm d}v)=0$ a.s. when $v\neq z$. This is because by Fubini’s theorem
$\mathbb{E}_{x}\left[\int_{\mathcal{I}}\tilde{p}(T_{z};x,v)m({\rm d}v)\right]$ $=\int_{\mathcal{I}}\mathbb{P}_{x}(X_{T_{z}}\in{\rm d}v)=\int_{\mathcal{I}}\int_{0}^{\infty}\mathbb{P}_{x}(X_{s}\in{\rm d}v,T_{z}\in{\rm d}s)$
$=\int_{0}^{\infty}\mathbb{P}_{x}(X_{s}\in{\rm d}z,T_{z}\in{\rm d}s)=\mathbb{P}_{x}(X_{T_{z}}\in{\rm d}z)=\mathbb{E}_{x}[\tilde{p}(T_{z};x,v)\delta_{z}({\rm d}v)]$
and $\mathbb{P}_{x}(X_{T_{z}}\in{\rm d}z)=1$ by the definition of $T_{z}$. Hence the random measure $\tilde{p}(T_{z};x,v)m({\rm d}v)$ has a point mass at $z$. With this preparation the second term becomes
@@
$\mathbb{E}_{x}\left[p(t-H_{z};X_{H_{z}},y)\mbox{1}\hskip-2.5pt\mbox{l}_{\{H_{z}<\lambda_{\alpha}<t\}}\right]$ $=\mathbb{E}_{x}\left[\int_{\mathcal{I}}\tilde{p}(T_{z}-H_{z};z,u)\tilde{p}(t-T_{z};u,y)m({\rm d}u)\mbox{1}\hskip-2.5pt\mbox{l}_{\{H_{z}<\lambda_{\alpha}<T_{z}<t\}}\right]$
$=\mathbb{E}_{x}\left[p(t-T_{z};X_{T_{z}},y)\mbox{1}\hskip-2.5pt\mbox{l}_{\{H_{z}<\lambda_{\alpha}<T_{z}<t\}}\right]$
@@
$=\mathbb{E}_{x}\left[\frac{h_{\alpha}(X_{T_{z}})}{h_{\alpha}(y)}p^{h_{\alpha}}(t-T_{z};X_{T_{z}},y)\mbox{1}\hskip-2.5pt\mbox{l}_{\{H_{z}<\lambda_{\alpha}<T_{z}<t\}}\right]$
@@
$=\mathbb{E}_{x}\left[\frac{1}{h_{\alpha}(y)}p^{h_{\alpha}}(t-T_{z};X_{T_{z}},y)\mbox{1}\hskip-2.5pt\mbox{l}_{\{H_{z}<\lambda_{\alpha}<T_{z}<t\}}\right],$
where $h_{\alpha}(X_{T_{z}})=1$ is due to the definition of $T_{z}$. Note that $\mbox{1}\hskip-2.5pt\mbox{l}_{\{\lambda_{\alpha}<H_{z}<t\}}+\mbox{1}\hskip-2.5pt\mbox{l}_{\{H_{z}<\lambda_{\alpha}<T_{z}<t\}}=\mbox{1}\hskip-2.5pt\mbox{l}_{\{T_{z}<t\}}$ by the assumption that $T_{z}<t$ whenever $\lambda_{\alpha}<t$. Combine these results to obtain
$\mathbb{P}_{x}(X_{t}\in{\rm d}y,\lambda_{\alpha}<t)=\mathbb{E}_{x}\left[p^{h_{\alpha}}(t-T_{z};X_{T_{z}},y)\mbox{1}\hskip-2.5pt\mbox{l}_{\{T_{z}<t\}}m({\rm d}y)\right]=\int_{0}^{t}p^{h_{\alpha}}(t-s;z,y)\mathbb{P}_{x}(T_{z}\in{\rm d}s)m({\rm d}y).$
@@
Since $z$ is arbitrary, we can make $z$ as close to $\alpha$ as we please. Therefore, for any time $\lambda_{\alpha}+u,u>0$, the post- last passage time process is governed by (3.4) for $u$ arbitrarily close to zero. ∎
@@
Although $h_{\alpha}(x)$ is not an excessive function, the transform (3.4) has the same form as if it were an excessive function. (See the literature listed in Section 1 for Doob’s $h$-transform.) Therefore, we can compute the infinitesimal drift and diffusion parameters of the transform (3.3) in the same way as if it is an $h$-transform. But there is a caveat that $x\neq\alpha$.
@@
Suppose that $X$ is a diffusion (does not have to be transient) and let $h$ be an excessive function. Let us denote the infinitesimal drift and diffusion parameters of the $h$-transform by $\mu^{h}(\cdot)$ and $\sigma^{h}(\cdot)$, respectively. The following result slightly generalizes the argument in Section 15.9 of Karlin and Taylor (1981).
@@
Assume $h$ is an excessive function and its derivative $h^{\prime}(y)$ exists for $y\in\mathcal{I}$. The infinitesimal drift $\mu^{h}(y)$ and diffusion parameter $\sigma^{h}(y)$ of the $h$-transform of $Y$ are
$\mu^{h}(y)=\mu(y)+\frac{h^{\prime}(y)}{h(y)}\sigma^{2}(y)\quad\text{and}\quad\sigma^{h}(y)=\sigma(y).$
@@
Turning to our case of $h_{\alpha}$,
$h_{\alpha}(x)=\mathbb{P}_{x}(\lambda_{\alpha}=0)=\frac{s(\alpha)-s(x)}{s(\alpha)-s(\ell)},\quad x\leq\alpha.$ (3.5)
@@
Before proceeding further, we should check if the point $\alpha$ is an entrance boundary for the post-last passage time process $(X_{t})_{t>\lambda_{\alpha}}$: the process can enter into $(-\infty,\alpha)$ from $\alpha$ but cannot exit from $\alpha$. This confirms that $(X_{t})_{t>\lambda_{\alpha}}$ cannot return to the region $\lambda_{\alpha},\infty)$.
@@
The point $\alpha$ is an entrance boundary for the post-last passage time process $X^{h_{\alpha}}_{t}(\omega):=\{\omega(t):\lambda_{\alpha}<t<H_{\ell},\mathbb{P}^{h_{\alpha}}_{\alpha}\}$ with $H_{\ell}:=\inf\{t:\omega(t)=\ell\}$.